OpenStudy (anonymous):

Help please!!! Find the derivative at each critical point and determine the local extrema values. f(x)=x^2(sqrt(3-x))

OpenStudy (cwrw238):

this is an equation not a function do you mean f(x) = x^2(sort(3 - x))

OpenStudy (anonymous):

Yes. sorry for the typo. @cwrw238

OpenStudy (anonymous):

fixed it!(:

OpenStudy (bahrom7893):

use the product and the chain rules to find the derivative: \[f'(x)=x^2*\frac{1}{2}(3-x)^{-\frac{1}{2}}*(-1)+2x\sqrt{3-x}\]

OpenStudy (anonymous):

once I derive it what do I do next? @bahrom7893

OpenStudy (loser66):

to me, I do something like \[f(x) = x^2 \sqrt{3-x}=\sqrt{3x^4-x^5}=(3x^4-x^5)^{\frac{1}{2}}\]take derivative this f(x) I get \[f'(x) = \dfrac{12x^3 -5x^4}{2\sqrt{3x^4-x^5}}\]then solve for x by letting numerator =0; combine with the condition of denominator not =0.

OpenStudy (anonymous):

so 12x^3-5x^4 =0? then solve for x?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

x=0 and x=12/5?

OpenStudy (loser66):

reject x =0 , because it make denominator =0 leads to the whole f'(x) undefined

OpenStudy (anonymous):

okay.. that makes sense,

OpenStudy (loser66):

the condition for denominator not =0 is 0<x<3 and your critical point x = 12/5 satisfies that condition. BINGOOOOOO! go ahead, friend

OpenStudy (anonymous):

Ohhhh... so how do i find if theres any local extrema values @Loser66

OpenStudy (loser66):

so far, you have 3 critical points x =0, x = 3 and x =12/5, right? plug them back to the original function to find out which f is biggest/ smallest . I mean f(0), f(3) and f (12/5)

OpenStudy (anonymous):

the biggest valuse would be the local max right?

OpenStudy (loser66):

yup